### v1

Abstract
We propose a novel framework that integrates medical image segmentation and point cloud upsampling for accurate shape reconstruction of pelvic models. Using the SAM-Med3D model for segmentation and a point cloud upsampling network trained on the MedShapeNet dataset, our method transforms sparse medical imaging data into high-resolution 3D bone models. This framework leverages prior knowledge of anatomical shapes, achieving smoother and more complete reconstructions. Quantitative evaluations using metrics such as Chamfer Distance etc, demonstrate the effectiveness of the point cloud upsampling in pelvic model. Our approach offers potential applications in reconstructing other skeletal structures, providing a robust solution for medical image analysis and statistical shape modeling.
Keywords:
Statistical Shape Modeling Point Cloud Upsampling Bone Shape Pelvic Segmentation.
1 Introduction
Figure 1: The flowchart of Point Cloud Upsampling as Statistical Shape Model proposed by us requires training a point cloud upsampling network to learn the distribution of bone shapes before inputting volumetric medical images. Subsequently, the image is input into the segmentation model, converted into a point cloud, and upsampled to reconstruct the complete model.
Statistical Shape Models (SSM) are based on traditional statistical methods designed to describe the average shape and density distribution, as well as the main variation patterns within these distributions. It is undeniable that U-Net[22] and its various deep learning segmentation models based on the U-Net architecture[7, 23] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[15, 34]. Additionally, various Segment Anything Model (SAM) approaches for medical imaging[12, 26, 18] are becoming increasingly prevalent. Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], automatic recognition[4], and medical diagnosis[11]. Patient-specific models based on SSM are used in numerous medical applications, such as generating patient-specific bone models from sparse 3D data or 2D image sets using 3D SSM[24], or for in-vivo bone motion tracking[3].
Creating ground truth for SSM differs from the masks used in semantic segmentation. Training an SSM requires manually identifying shape landmarks to infer statistical data. To build these shapes, a template grid is gradually deformed to segment training images and match the target boundaries, such as those of organs or bones.
To evaluate new patient scans using SSM, the new scan must be segmented and aligned, and the SSM must be optimized for the entire shape population. In this regard, Adams and Elhabian[1] introduced an unsupervised method for predicting the Point Distribution Model (PDM), enabling point cloud networks to learn accurate anatomical SSMs. Moreover, deep learning methods, inspired by renowned shape benchmarks in computer vision (CV) such as ShapeNet[6], have further developed MedShapeNet[14], which is based on real patient imaging data and instruments.
Building on this, our work focuses on the Pelvic1k dataset[16] with the goal of learning realistic bone shape features through point cloud upsampling. We sample the Pelvic model from MedShapeNet as a point cloud, allowing the upsampling model to learn these features. We then construct a framework that employs the SAM-Med3D model to segment medical images, converts the segmented voxel results into point clouds, and applies the upsampling model, which has prior knowledge of real data, to perform point cloud upsampling. This process transforms original, resolution-limited medical images and sparse voxels into denser, high-resolution 3D bone models, thereby improving shape reconstruction accuracy. The integration of medical image segmentation and point cloud upsampling within this framework enables SSM to describe the average shape and facilitates segmentation and model reconstruction.
In summary, the key contributions of our work are as follows:
1) We demonstrate the necessity of learning prior knowledge from MedShapeNet by sampling the pelvic model as point clouds and constructing a baseline using two SOTA point cloud upsampling models, comparing the results with those based on PU1k.
2) We apply the SAM-Med3D model to the Pelvic1k dataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model. This process results in smoother, more complete 3D shape reconstructions.
3) The proposed framework, which integrates medical image segmentation and point cloud upsampling, has the potential to be transferred to various other skeletal models and organs.
2 Related Works
2.1 Point Cloud Processing
Since the introduction of PointNet[19, 20], this direct processing avoids the loss of context and complex steps involved in the data conversion process. Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details. With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks. For example, in point cloud processing tasks, PCN[31] first employs a similar encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input. Later, ASFM-Net[27] used parallel global and local feature matching to reasonably infer missing geometric details of objects.
Besides, PU-Net[30] is a pioneering work that introduces CNNs for point cloud upsampling. Subsequently, Grad-PU[10] similarly provides flexible upsampling at different scales. PUCRN[8] optimized point cloud generation and performance through a multi-stage structure and Transformers[25, 33], especially in sparse or missing regions.
In medical image applications, point cloud completion was first explored by Adams and Elhabian[1], who used [31] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds. And Gu et al. [9] proposed a new approach for 3D reconstruction of individual X-ray images, using point cloud completion to reconstruct real bones from unseen CT images. Later, Yassin et al.[28] proposed a multimodal foundation model for medical point cloud completion based on MedShapeNet[14]. This model is designed to handle incomplete 3D point cloud data and reconstruct the complete shape of various medical structures. By combining 3D point cloud data with textual data, the model improves the accuracy of shape reconstruction. Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
2.2 Segment Anything
The Segment Anything Model (SAM)[12] was trained on a dataset of 11 million images and over 1 billion masks. This design allows the model to be enabled zero-shot transfer to new image distributions and tasks. Due to its speed and versatility, SAM’s zero-shot capabilities have led to its use in medical image segmentation tasks, giving rise to MedSAM[18]. The MedSAM model[18] was developed on a large-scale medical image dataset containing 1.5 million images and masks, covering various imaging modalities and tumor types. It offers better accuracy and robustness than modality-specific models. However, MedSAM[18] only trained by organs and skin, including endoscopic images, and lacks 3D medical images. Later, SAM-Med3D[26] was introduced, focusing on volumetric medical datasets for training and producing satisfactory 3D results. Notably, its training targets include not only organs, skin, and cells but also bones. It is worth mentioning that SSM represents shape variation through statistical analysis of anatomical variations in shape populations generated from medical images, and it has applications in multiple medical fields, including facial recognition and skull analysis[5]. This drives our objective of combining deep learning methods for medical image segmentation and point cloud upsampling, enabling SSM modeling and shape variation.
3 Methodology
3.1 Proposed Framework
In this paper, we propose a framework called Point Cloud Upsampling as Statistical Shape Model (PUSSM), which combines deep learning-based medical image segmentation with point cloud upsampling. Specifically, we first use a real medical shape dataset based on MedShapeNet and apply it to a point cloud upsampling network. This enables us to learn prior knowledge of real-world medical shapes, allowing the upsampling of point clouds from sparse to dense points in anatomical shapes, as shown in Figure 1 (a).
Next, as illustrated in Figure 1 (b), volumetric medical images are input into the pre-trained SAM-Med3D model to obtain segmentation voxel results, which are then converted into point clouds using spatial coordinates. These sparse point clouds are then fed into the point cloud upsampling network to generate a more refined point cloud. Finally, surface reconstruction is performed on the refined point cloud to obtain a denser, higher-resolution volumetric medical model.
The entire process follows the concept of Statistical Shape Models (SSM)[13], where accurate anatomical shapes are learned from the dataset (prior knowledge), and segmentation is used for modeling. Ultimately, similar to visual methods, we achieve segmentation and modeling through SSM’s description of the average shape. However, our network is not end-to-end; instead, it consists of two separate networks: one for medical image segmentation and the other for point cloud upsampling.
3.2 Evaluation Metrics
To evaluate our point cloud upsampling network, we used three key metrics: Chamfer Distance (CD), Hausdorff Distance (HD), and Point-to-Surface Distance (P2F).
Chamfer Distance (CD) measures the average of the nearest-neighbor distances between two point sets.
Hausdorff Distance (HD) captures the maximum distance between points in one set and their closest point in the other set, thus highlighting the worst-case mismatch.
Point-to-Surface Distance (P2F) evaluates the shortest distance from upsampled points to the surface of the reference model.
4 Experiments
4.1 Settings
4.1.1 Datasets
In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[14] for point cloud upsampling to learn shape prior knowledge. We then visualized subject’s result based on Pelvic1k[16], a comprehensive pelvic CT dataset capable of replicating actual appearance changes.
For training and testing the upsampling network, we used the MedShape Pelvic dataset based on MedShapeNet[14], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[29]. The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples. Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
4.1.2 Comparison Methods
To verify the effectiveness of our proposed theory in point cloud upsampling as statistical shape model, comparing it with two advanced algorithms: Grad-Pu[10] that realized any ratio upsampling and PUCRN[8] that is the SOTA model.
For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment. And we test the result based on training in MedShape Pelvic[14] and PU1k dataset[21], which is a synthetic dataset from 3D models. This comparison with identical parameter settings would be presented in ablation: Section 4.3.
4.1.3 Implementation Details
Experiments were developed using the PyTorch framework on the Ubuntu 22.04 system. We used an NVIDIA L20 GPU with 48GB of graphics memory and an Intel(R) Xeon(R) Platinum 8457C host with 100GB of RAM. The network was trained over 100 epochs with a batch size of 64. The initial learning rate was set to 0.001, with a decay rate of 0.05.
4.2 Upsampling Results
In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShape Pelvic dataset in 4×4$\times$4 × upsampling. The quantitative results for the dataset in Table 1 show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
This result demonstrates the effectiveness of PUCRN in CD, HD, and P2F, showcasing its superior capability in learning the shape prior knowledge in pelvic. The 3 stage network refinement structure allow it could minimizing nearest-neighbor distance differences and reducing maximum point-to-point distance errors.
Additionally, noticed that Grad-PU with very high HD. This may be caused by the input pelvic training data. This reflects that Grad PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 2 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
Besides, we also conducted 16×16$\times$16 × upsampling tests, in Table 2. Grad PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set. However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two 4×4$\times$4 × samplings on PUCRN to achieve a 16×16$\times$16 × sampling effect. The results show that PUCRN still has good performance.
Table 1: Performance comparison of two methods with MedShape Pelvic across input scales. Among them, the unit is $10^{-3}$.
Methods 512 input 1024 input 2048 input 4096 input
[…]
Grad-PU 5.643 124.639 12.438 3.682 82.969 9.779 2.477 49.451 4.852 1.690 28.871 3.172
PUCRN 3.104 7.806 6.150 3.946 16.635 6.436 1.839 5.709 4.066 1.274 5.087 2.719
4.3 Ablation Study
To verify the effectiveness and necessity of training in MedShape Pelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShape Pelvic[14] and PU1k dataset[21] have been tested. Among them, MedShape Pelvic has 50496 training samples and PU1k has 69000 training samples. This avoids the difference in learning effectiveness caused by the large difference in the number of training samples. The result shows in Table 3
Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024. Among them, the unit is $10^{-3}$.
Methods 512 to 8192 1024 to 16384
[…]
Grad-PU 5.827 124.640 13.117 3.422 84.825 10.329
PUCRN 3.586 9.984 6.824 2.870 18.899 5.872
As a result, training on the MedShape Pelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features. This is very important. However, Grad PU’s model on the PU1k dataset performed better than MedShape Pelvic’s test set. These reasons are multifaceted, and Grad PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShape Pelvic obtained from real volume medical images.
Table 3: Pelvic point cloud upsampling results in two different train sets.
Methods 512 input 1024 input 2048 input 4096 input
[…]
Grad-PU(MedShapeNet) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
PUCRN(MedShapeNet) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
Grad-PU(PU1k) 5.411 15.551 3.353 12.280 2.256 8.677 1.513 7.872
PUCRN(PU1k) 3.749 10.065 2.902 9.955 1.995 6.134 1.379 5.253
4.4 Visualization
In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM. In this visualization section, due to the lack of ground truth, we can only choose to display the reconstruction effect. If we want to make quantitative comparisons, it involves the need for No Reference Mesh Quality Assessment. However, mostly 3D quality assessment focusing on point cloud instead of mesh [32].
Figure 2: Point cloud upsampling results on Pelvic1k segment sparse inputs. Compared with the voxel surface in segmented results, our method can generate more uniform with detailed structures.
Now we input the Pelvic1k dataset into SAM Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[17]. As shows in Figure 2, compared with the surface obtained by directly segmenting voxels through marching cubes (Pseudo Ground Truth), our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN.
5 Conclusion
This study introduced a framework integrating point cloud upsampling with statistical shape modeling to improve high-resolution pelvic model reconstruction. By combining SAM-Med3D segmentation with upsampling networks, we transformed sparse voxel data into dense point clouds and reconstructed detailed surfaces using the Marching Cube algorithm. Experiments demonstrated the necessity of upsampling and confirmed the framework’s effectiveness in achieving smoother and more complete reconstructions.
Future work will extend this framework to other anatomical structures and incorporate no reference quality assessment for objective evaluations.
{credits}
5.0.1 Acknowledgements
This research was supported by the National Natural Science Foundation of China under Grant 61773164.
[…]
The Department of Rehabilitation Science, The Hong Kong Polytechnic University, has not financial support, and there are no potential conflicts of interest to disclose.