### v3

# PUSSM: Point Cloud Upsampling as Implicit Statistical Shape Model

Abstract
This paper proposes a framework for high-fidelity reconstruction of pelvic structures by integrating medical image segmentation and point cloud upsampling. By point cloud upsampling to learn shape priors from MedShapePelvic without requiring landmarks or PCA, our method functions as an implicit statistical shape model. Evaluations on Pelvic1k show significant improvements in surface quality and anatomical accuracy. This approach is generalizable and applicable to other skeletal regions.
Keywords:
Statistical Shape Modeling Point Cloud Upsampling Bone Shape Pelvic Segmentation.
1 Introduction
Statistical Shape Models (SSM) are based on traditional statistical methods designed to describe the average shape and density distribution, as well as the main variation patterns within these distributions. It is undeniable that U-Net[25] and its various deep learning segmentation models based on the U-Net architecture[6, 26] have consistently achieved state-of-the-art (SOTA) results in organ and bone segmentation tasks[36]. Additionally, various Segment Anything Model (SAM) approaches for medical imaging[15, 28, 21] are becoming increasingly prevalent. Despite this, SSM remains widely used across various domains, particularly in tasks such as image segmentation and measurement[2], and medical diagnosis[14]. Patient-specific models based on SSM are used in numerous medical applications, such as generating patient-specific bone models from sparse 3D data or 2D image sets using 3D SSM[27], or for in-vivo bone motion tracking[3].
Figure 1: The proposed Point Cloud Upsampling as Statistical Shape Model involves training a point cloud upsampling network to learn bone shape distributions from volumetric medical images. The image is then segmented into a point cloud, upsampled, and reconstructed into a complete model.
Creating ground truth for SSM differs from the masks used in semantic segmentation. Training an SSM requires manually identifying shape landmarks to infer statistical data. To build these shapes, a template grid is gradually deformed to segment training images and match the target boundaries, such as those of organs or bones.
To evaluate new patient scans using SSM, the new scan must be segmented and aligned, and the SSM must be optimized for the entire shape population. In this regard, Adams and Elhabian[1] introduced an unsupervised method for predicting the Point Distribution Model (PDM), enabling point cloud networks to learn accurate anatomical SSMs. Moreover, deep learning methods, inspired by renowned shape benchmarks in Computer Vision (CV) have further developed MedShapeNet[16], which is based on real patient imaging data and instruments. However, classical SSM construction typically involves manual annotation of anatomical landmarks and tedious correspondence alignment, which may limit scalability and automation in large datasets.
In contrast, our approach explores a data-driven alternative to SSM, where prior anatomical knowledge is learned implicitly through point cloud upsampling networks trained on large-scale 3D medical datasets, which is on the Pelvic model of MedShapeNet dataset (MedShapePelvic)[16] with the goal of learning realistic bone shape features through point cloud upsampling. Instead of modeling shape variation explicitly via PCA, our method leverages learned feature distributions from dense anatomical shapes to enhance the reconstruction of sparse pelvic scans. This shift from explicit modeling to learned priors enables more flexible and automatic shape analysis, especially when integrated with modern segmentation tools such as SAM-Med3D. This process transforms original, resolution-limited medical images and sparse voxels into denser, high-resolution 3D bone models, thereby improving shape reconstruction accuracy. The integration of medical image segmentation and point cloud upsampling within this framework enables SSM to describe the average shape and facilitates segmentation and model reconstruction.
In summary, the key contributions of our work are as follows:
1) We demonstrate the necessity of learning prior knowledge from MedShapePelvic by sampling the pelvic model as point clouds and constructing a baseline using four SOTA point cloud upsampling models, comparing the results with those based on PU1k.
2) We apply the SAM-Med3D model to the Pelvic1k subdataset to obtain predicted segmented voxels, convert the voxels into point clouds, and input them into the upsampling model. This process results in smoother, more complete 3D shape reconstructions.
3) The proposed framework, which integrates medical image segmentation and point cloud upsampling, has the potential to be transferred to various other skeletal models and organs.
2 Related Works
2.1 Point Cloud Processing
2.1.1 In 3D Object
Since the introduction of PointNet[22, 23], this direct processing avoids the loss of context and complex steps involved in the data conversion process. Point cloud completion, in particular, is used to extract global features of 3D model, which can help supplement geometric details[18, 17]. With the aid of deep learning and large-scale 3D datasets, learning-based methods have achieved excellent performance in shape completion tasks. For example, in point cloud processing tasks, PCN[33] first employs a similar encoder to extract features and outputs dense and completed point clouds from sparse and incomplete input. Later, ASFM-Net[29] used parallel global and local feature matching to reasonably infer missing geometric details of objects.
Besides, PU-Net[32] is a pioneering work that introduces CNNs for point cloud upsampling. Subsequently, PU-GCN[24] employed a graph-based network construction, achieving good upsampling performance. Grad-PU[13] similarly provides flexible upsampling at different scales. PUCRN[9] optimized point cloud generation and performance through a multi-stage structure and Transformers[35], especially in sparse or missing regions. However, these methods are often trained on synthetic datasets (e.g., PU1k dataset[24] proposed by PU-GCN) and may struggle to generalize to real anatomical shapes with irregular geometry and sparse sampling. Moreover, they typically focus on object-level CAD models and are not directly validated on medical segmentation pipelines.
2.1.2 In Medical Image
For medical image applications, Chen et al.[7] proposed a learning grid to point cloud deformable shape registration framework from multiple 2D image slices, which can achieve accurate 3D grid reconstruction in real-time in the absence of data and sparse annotations. Point cloud completion was first explored by Adams and Elhabian[1], who used [33] in an unsupervised manner to predict the Point Distribution Model (PDM) from 3D point clouds, exploring the potential of learning SSM from point clouds. And Gu et al. [10] proposed a new approach for 3D reconstruction of individual X-ray images, using point cloud completion to reconstruct real bones from unseen CT images. Later, Yassin et al.[30] proposed a multimodal foundation model for medical point cloud completion based on MedShapeNet[16]. This model is designed to handle incomplete 3D point cloud data and reconstruct the complete shape of various medical structures. By combining 3D point cloud data with textual data, the model improves the accuracy of shape reconstruction. Zhang et al.[34] proposed a hierarchical feature learning framework based on the State Space Model for medical point cloud understanding, which will aid in anatomical classification, completion, and segmentation.Although point cloud completion shares similarities with point cloud upsampling tasks by extracting global features from partial input to complement geometric details, no current work has applied these techniques to 3D point cloud data in the context of SSM to reconstruct various medical structures.
In contrast to these works, our approach explicitly integrates a medical image segmentation model (SAM-Med3D) with point cloud upsampling to form a complete pipeline from imaging to shape modeling. Furthermore, by training on realistic datasets, MedShapePelvic, and test on Pelvic1k subdataset, our method captures clinically relevant anatomical features and improves the quality of reconstructed surfaces.
2.2 Segment Anything
The Segment Anything Model (SAM)[15] was trained on a dataset of 11 million images and over 1 billion masks. This design allows the model to be enabled zero-shot transfer to new image distributions and tasks. Due to its speed and versatility, SAM’s zero-shot capabilities have led to its use in medical image segmentation tasks, giving rise to MedSAM[21]. The MedSAM model[21] was developed on a large-scale medical image dataset containing 1.5 million images and masks, covering various imaging modalities and tumor types. It offers better accuracy and robustness than modality-specific models. However, MedSAM[21] only trained by organs and skin, including endoscopic images, and lacks 3D medical images, because they focus on interactive in 2D images. Later, SAM-Med3D[28] was introduced, focusing on volumetric medical datasets for training and producing satisfactory 3D results. Notably, its training targets include not only organs, skin, and cells but also bones. It is worth mentioning that SSM represents shape variation through statistical analysis of anatomical variations in shape populations generated from medical images, and it has applications in multiple medical fields, including facial recognition and skull analysis[5]. This drives our objective of combining deep learning methods for medical image segmentation and point cloud upsampling, enabling SSM modeling and shape variation.
Ergo, our study leverages SAM-Med3D as a frontend for downstream statistical shape modeling. By combining its segmentation output with point cloud upsampling, we demonstrate that deep segmentation models can serve as a viable input source for high-fidelity shape reconstruction.
3 Methodology
3.1 Overview of the Proposed Framework
We propose a novel framework called Point Cloud Upsampling as Implicit Statistical Shape Model (PUSSM), which integrates medical image segmentation and point cloud upsampling for high-fidelity anatomical reconstruction. As shown in Figure 1(a), the core idea is to learn an implicit shape prior from real anatomical data through a deep upsampling network, allowing sparse point clouds to be restored to dense representations in an anatomically plausible way.
The pipeline consists of two sequential modules. First, we apply the SAM-Med3D model to volumetric CT data to obtain voxel-wise segmentation masks. These masks are converted into sparse point clouds using spatial coordinates. Then, as illustrated in Figure 1(b), the sparse point cloud is passed into a pretrained point cloud upsampling network, which has been trained on MedShapePelvic samples. Finally, surface reconstruction is applied to the dense point cloud using Marching Cubes.
Figure 2: Distributional interpretation of implicit statistical shape modeling. The segmentation-derived point cloud (left) deviates from the real anatomical distribution. Our network $f_{\theta}$ learns to map the sparse inputs toward the ground-truth distribution through point cloud upsampling.
3.2 Implicit Statistical Shape Modeling
Unlike classical Statistical Shape Models (SSMs), which explicitly encode anatomical variation via PCA over aligned meshes, our framework learns an implicit statistical shape model through data-driven training. Specifically, let $X\subset\mathbb{R}^{3}$ denote a sparse point cloud from SAM-Med3D, and let $Y\subset\mathbb{R}^{3}$ be the corresponding ground truth dense point cloud from MedShapePelvic.
We train a deep function $f_{\theta}:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{M\times 3}$ to minimize the geometric discrepancy between $f_{\theta}(X)$ and $Y$, using a distance metric $D(\cdot,\cdot)$ such as the Chamfer Distance:
$\min_{\theta}\mathbb{E}_{(X,Y)\sim\mathcal{D}}\left[D(f_{\theta}(X),Y)\right]$
This process embeds the shape prior directly into the network weights $\theta$, enabling the model to generalize to new, noisy, or incomplete segmentation results. Thus, $f_{\theta}$ serves as an implicit statistical shape model learned directly from anatomical data—requiring no landmark alignment, shape vectors, or PCA decomposition.
Figure 2 further illustrates this concept from a distributional perspective, showing how our model maps the segmentation-derived distribution toward the ground-truth anatomical distribution via upsampling.
Figure 3: Conceptual comparison between classical SSM and our implicit PUSSM framework across key dimensions.
3.3 Comparison with Classical SSMs
To highlight the differences between our method and classical SSMs, Figure 3 provides a conceptual comparison across three key dimensions: shape representation, prior data source, and learning strategy. Traditional SSMs rely on manually constructed shape vectors and PCA decomposition of landmark-aligned meshes. In contrast, our method encodes shape priors directly through end-to-end learning from dense mesh data, enabling greater flexibility and broader applicability.
3.4 Evaluation Metrics
3.4.1 For Point Cloud Upsampling
To evaluate the point cloud upsampling network, we used three key metrics:
Chamfer Distance (CD) measures the average of the nearest-neighbor distances between two point sets. Let $P$ and $Q$ be the two sets of points (the upsampled point cloud and the ground truth point cloud, respectively), the Chamfer Distance is defined as:
$\text{CD}(P,Q)=\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|p-q\|^{2}+\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|q-p\|^{2}$
Hausdorff Distance (HD) captures the maximum distance between points in one set and their closest point in the other set, thus highlighting the worst-case mismatch. The Hausdorff distance between sets $P$ and $Q$ is defined as:
$\text{HD}(P,Q)=\max\left(\max_{p\in P}\min_{q\in Q}\|p-q\|,\max_{q\in Q}\min_{p\in P}\|q-p\|\right)$
Point-to-Surface Distance (P2F) evaluates the shortest distance from upsampled points to the surface of the reference model. Given a set of upsampled points $P$ and a reference surface $S$, the Point-to-Surface Distance is defined as:
[…]
3.4.2 For Surface Reconstruction
To evaluate the surface reconstruction of our framework, we used three key metrics: Area-Length Ratio, Manifoldness Rate and Connected Component Discrepancy[11, 4].
Area-Length Ratio (ALR) measures the shape quality of individual triangles within a mesh. This ratio attains its maximum value of 1 for equilateral triangles, indicating optimal shape quality. It is defined as:
Manifoldness Rate (MR) quantifies the proportion of edges in a mesh that satisfy manifold conditions. An edge is considered manifold if it is shared by exactly two faces. A rate approaching 1 indicates a well-structured mesh suitable for operations like Boolean computations and 3D printing. Conversely, a lower rate highlights the presence of non-manifold edges, which may necessitate mesh repair or refinement. It is defined as:
Manifoldness Rate=Number of manifold edgesTotal number of $\text{Manifoldness Rate}=\frac{\text{Number of manifold edges}}{\text{Total number of edges}}$
Connected Component Discrepancy (CC Diff.) assesses the topological connectedness of a mesh by comparing the observed number of connected components to the expected number. A discrepancy of 0 signifies that the mesh has the anticipated number of connected components, implying structural integrity. It is defined as:
[…]
Those three indicators to evaluate the geometric quality of reconstructed meshes.
Besides, to evaluate the fidelity of the reconstructed meshes against the ground truth, we employ several widely recognized metrics[8]: Chamfer Distance (CD), F-Score (F1), Normal Consistency (NC), Edge Chamfer Distance (ECD), and Edge F-Score (EF1).
4 Experiments
4.1 Settings
4.1.1 Datasets
In our study, experimental validation on the pelvic region have been conducted in the hips, femurs, and sacrum from MedShapeNet[16] as MedShapePlevic for point cloud upsampling to learn shape prior knowledge. We then visualized Quantitative analyzed result based on Pelvic1k[19], a comprehensive pelvic CT dataset capable of replicating actual appearance changes. Due to the lack of maintenance in some of the open source data used by Pelvic1k, we can only obtain the following data (e.g. subdatasets 3, 4, 6).
For training and testing the upsampling network, we used the MedShapePelvic dataset based on MedShapeNet[16], which includes 50,496 training samples from 3,156 Mesh models generated via patch sampling[31]. The dataset consists of 3,506 3D models, split into 3,156 training samples and 350 testing samples. Patch sampling was performed by generating training data through Poisson disk sampling on patches of the 3D meshes.
4.1.2 Comparison Methods
In order to verify the effectiveness of our proposed theory, point cloud upsampling is the first part. We built a baseline based on MPU[31], PU-GCN[24], Grad-Pu[13] and PUCRN[9]. In more experiments, only by comparing it with two advanced algorithms: Grad-Pu that realized any ratio upsampling and PUCRN that is the SOTA model.
For a fair and objective comparison, we obtained the open-source codes of these methods and trained and tested them on our computing equipment. And we test the result based on training in MedShapePelvic[16] and PU1k dataset[24], which is a synthetic dataset from 3D models. This comparison with identical parameter settings would be presented in ablation: Section 4.3.
4.1.3 Implementation Details
Experiments were developed using the PyTorch framework on the Ubuntu 22.04 system. We used an NVIDIA L20 GPU with 48GB of graphics memory and an Intel(R) Xeon(R) Platinum 8457C host with 100GB of RAM. The network was trained over 100 epochs with a batch size of 64. The initial learning rate was set to 0.001, with a decay rate of 0.05.
4.2 Upsampling Results
In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShapePelvic dataset in 4×4$\times$4 × upsampling. The quantitative results for the dataset in Table 1 show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
This result demonstrates the effectiveness of PUCRN in CD, HD, and P2F, showcasing its superior capability in learning the shape prior knowledge in pelvic. The 3 stage network refinement structure allow it could minimizing nearest-neighbor distance differences and reducing maximum point-to-point distance errors.
Additionally, noticed that Grad-PU with very high HD. This may be caused by the input pelvic training data. This reflects that Grad-PU, lacking local and global refinement, cannot achieve geometric differences based on Pelvic training data, in Figure 4 showed that the upsampling results of Grad-PU have many holes which means it cannot learn good enough in local features.
Besides, we also conducted 16×16$\times$16 × upsampling tests, in Table 2. Grad-PU can perform upsampling at any scale, so we set its ratio directly to 16×16$\times$16 × on the test set. However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two 4×4$\times$4 × samplings on PUCRN to achieve a 16×16$\times$16 × sampling effect. The results show that PUCRN still has good performance.
Table 1: Performance comparison of two methods with MedShapePelvic across input scales.
Method 512 input 1024 input 2048 input 4096 input
[…]
MPU 13.050 11.971 10.604 8.507 26.088 8.759 5.718 9.421 8.558 3.906 8.768 7.556
PU-GCN 4.191 9.942 8.553 5.326 21.284 7.329 2.483 7.296 4.652 1.719 6.519 3.671
Grad-PU 5.643 124.639 12.438 3.682 82.969 9.779 2.477 49.451 4.852 1.690 28.871 3.172
PUCRN 3.104 7.806 6.150 3.946 16.635 6.436 1.839 5.709 4.066 1.274 5.087 2.719
4.3 Ablation Studies
To verify the effectiveness and necessity of training in MedShapePelvic to learn the prior knowledge of shape, the 4×4$\times$4 × sampling result based on training in MedShapePelvic and PU1k dataset[24] have been tested. Among them, MedShapePelvic has 50496 training samples and PU1k has 69000 training samples. This avoids the difference in learning effectiveness caused by the large difference in the number of training samples. The result shows in Table 3
Table 2: Quantitative comparison in 16×16$\times$16 × upsampling after training in 4×4$\times$4 ×. The input points are 512 and 1024.
Method 512 to 8192 1024 to 16384
[…]
Grad-PU 5.827 124.640 13.117 3.422 84.825 10.329
PUCRN 3.586 9.984 6.824 2.870 18.899 5.872
As a result, training on the MedShapePelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features. This is very important. However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set. These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
Table 3: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
Method 512 input 1024 input 2048 input 4096 input
[…]
Grad-PU(MedShapePelvic) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
PUCRN(MedShapePelvic) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
Grad-PU(PU1k) 5.411 15.551 3.353 12.280 2.256 8.677 1.513 7.872
PUCRN(PU1k) 3.749 10.065 2.902 9.955 1.995 6.134 1.379 5.253
Figure 4: Point cloud upsampling results on Pelvic1k segment sparse inputs. Compared with the voxel surface in segmented results, our method can generate more uniform with detailed structures.
4.4 Reconstruction Visualizations
In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM.
The Pelvic1k subdataset 3,4 and 6 have been inputted into SAM-Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[20], where voxel size is $1.5$. As shows in Figure 4, compared with the surface of ground truth, and obtained by directly segmenting voxels through marching cubes. Our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN. And there are no obvious contours of layer after reconstructed.
4.5 Reconstruction Evaluations
To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in Section3.4.2 For Surface Reconstruction.
Table 4: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
Subdataset 3 Subdataset 4 Subdataset 6
[…]
Ground Truth 0.692 100.0 0.26 0.721 100.0 1.68 0.742 100.0 0.66
Segmented Result (by SAM-Med3D) 0.676 100.0 1.94 0.699 100.0 4.89 0.735 100.0 1.18
PUSSM (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
PUSSM (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
As shows in Table 4. Our method outperforms ground truth and segmented result in terms of mesh quality and achieves stable reconstruction accuracy. Specifically, our method has uniform grid partitioning (as shown in ALR) and eliminates floating artifacts (as shown in CC Diff.). This improvement is attributed to the finer points and distribution brought about by point cloud upsampling.
Furthermore, we conducted quantitative comparisons on the reconstruction results in order to evaluate surface reconstruction accuracy and feature preservation. We used metrics (CD, F1, NC, ECD, EF1) in Section3.4.2 For Surface Reconstruction.
To ensure a fair and uniform comparison, we employ Poisson disk sampling to generate 16,384 points from each mesh, including the ground truth, segmented results, and upsampled reconstructions. This sampling strategy maintains uniform point distribution and adequately captures both global structures and fine details, facilitating reliable metric computation.
Table 5: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
Subdataset 3 Subdataset 4 Subdataset 6
[…]
Segmented Result (by SAM-Med3D) 8.379 0.214 0.691 7.420 0.408 56.272 0.126 0.711 54.559 0.250 13.997 0.214 0.716 10.724 0.357
PUSSM (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
PUSSM (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
As shows in Table 5. Our method, across all three subdatasets, PUCRN consistently outperforms both Grad-PU and the raw SAM-Med3D segmentations. For example, on Subdataset 3, PUCRN achieves a significantly lower CD compared to Grad-PU, along with superior edge preservation as reflected by a lower ECD and a higher EF1-score. Similar trends are observed on Subdataset 4 and Subdataset 6, where PUCRN achieves the best overall performance in terms of both global surface quality and local edge fidelity. These results demonstrate the effectiveness of PUCRN in enhancing the downstream surface reconstruction quality by providing denser and more accurate point distributions.
5 Conclusion
In this work, we proposed PUSSM, an implicit statistical shape modeling framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data. Our method avoids the limitations of classical SSMs by eliminating the need for landmark annotations and explicitly modeling shape variation. Experiments demonstrate improved surface quality and anatomical plausibility across multiple evaluation metrics.
Future work will extend this framework to 3D SAM model[15] by points based prompting to integrate it as Chain-of-Thought (CoT)[12] for Segmentation.
{credits}